OpenStudy (anonymous):

The sum of four consecutive even integers is 308. Write and solve an euation to find the four integers

OpenStudy (saifoo.khan):

Suppose the number be x. Now, First # = x Second # = x+1 Third # = x + 2 Forth = x+3

OpenStudy (saifoo.khan):

Now, x + (x+1) + (x+2) + (x+3) = 308

OpenStudy (saifoo.khan):

Solve for x.

OpenStudy (anonymous):

Since the consecutive numbers are even, shouldn't it be: x + x + 2 + x + 4 + x + 6 = 308

OpenStudy (anonymous):

\[n+n+2+n+4+n+6=308\] \[4n+12=308\] \[4n=296\]etc